# Advanced Video Inpainting Using Optical Flow-Guided Efficient Diffusion

### v2

@@ 5 Experiments
FloED is built upon Stable Diffusion (SD) [1] and incorporates motion modules initialized from AnimateDiff v3 [7]. We employ RAFT [24] for flow estimation and SD v2 inpainting [2] for the anchor frame strategy.
For training, we utilize the video dataset from Open-Sora-Plan [14], which comprises over 400K high-quality videos sourced from Pexels and Pixabay platforms. During the training phase, we use 16 frames at a resolution of 256×256, and synthetic random masks are applied throughout the training process.
For the two-stage training process, the number of epochs is set to 5 and 30, respectively. And $\lambda$ is set to 0.1 during the second stage. A batch size of 128 is achieved through gradient accumulation across eight NVIDIA A800 GPUs.
During inference, we empirically define the speed-up step $S$ as 5 (25 steps). We have developed an evaluation benchmark comprising 100 previously unseen videos, with 50 designated for object removal (OR) and 50 for background restoration (BR). For BR task, we use synthetic random masks that focus on the background. For the OR task, object masks are obtained by applying Segment-Anything [12] (SAM) to each frame.
5.2 Qualitative Results
To comprehensively demonstrate the capabilities of our method, we evaluated FloED on real videos for object removal and background restoration tasks. As illustrated in Fig. 4, FloED effectively handles both OR and BR tasks across a wide range of mask sizes, including those covering more than half of the frame area. Our approach adeptly in-paints the specified regions with coherent and contextually appropriate content. Additionally, FloED exhibits excellent temporal consistency throughout the video.
@@ 5.3 Comparisons
We present a comprehensive evaluation of our method compared to other open-source video inpainting solutions, including VideoComposer [26], CoCoCo [38], and Propainter [37]. Building upon the same image inpainting [2] results from anchor frames, we also assess the performance of the representative I2V model Dynamicrafter [29] as another comparison. We generate the corresponding video prompts for BR using VideoGPT [18], while manually creating appropriate background prompts for OR. Following the prior method [37], we applied a copy-paste approach to replicate the out-of-mask region across all methods except Dynamicrafter [29].
@@
Fig. 5 (left) compares the performance of background restoration, while Fig. 5 (right) illustrates object removal. Compared to the representative transformer-based method Propainter [37], our approach generates more vivid and detailed results in object removal (right), whereas Propainter encounters difficulties when synthesizing new content (left). Regarding text-guided diffusion methods, CoCoCo [38] and VideoComposer [26] tend to produce hallucinated content that is inconsistent with the surrounding environment. When focusing on the shadow regions within the mask area (right), our method demonstrates superior temporal consistency compared to Dynamicrafter [29].
Task BR OR
Metrics PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑ TA ↑
DC 17.50 0.4157 0.934 6.17 0.985 22.44
VC 22.81 0.8614 0.193 3.43 0.987 21.30
Propainter 27.02 0.9057 0.129 1.51 0.996 20.94
CoCoCo 23.08 0.8694 0.165 3.73 0.991 21.97
Ours 29.17 0.9441 0.118 2.83 0.994 22.49
Table 1: Quantitative comparisons using different metrics, including PSNR, SSIM, FVID, Optical flow warping error, Temporal consistency (TC), and Text alignment (TA).
Figure 6: We conduct reliable User Study with randomized order to assess their performance across different inpaintng outcomes.
Quantitative comparisons.
(a) Metric Evaluation. For background restoration, we employ PSNR [28], VFID [25], and SSIM [27] to quantify basic quality. Additionally, we assess temporal consistency using flow warping error [6, 13] in conjunction with Temporal Consistency (TC) [36]. TC is measured by the cosine similarity between consecutive frames in the CLIP-Image [8] feature space. For object removal, since ground truth data is unavailable for evaluating the aforementioned metrics, we utilize Text Alignment (TA) [8] as an evaluation metric, leveraging the CLIP score. For consistency, all metric evaluations were conducted at a resolution of 512×512. As shown in Tab. 1, our method outperforms other methods in PSNR, SSIM, VFID, and TA. Additionally, it surpasses other diffusion-based methods in flow warp error and TC, second only to Propainter by a small margin, due to the inherent generative characteristics of diffusion. Overall, FloED delivers excellent temporal consistency and high-quality outcomes compared to the state-of-the-art solutions. (b) User study. Since CLIP scores do not always align with human perception [36], we conducted a comprehensive user study. 12 annotators evaluated different inpainting results of 100 videos from 5 methods, assessing temporal consistency, content coherence, and overall quality. As illustrated in Fig. 6, our model was highly favored, achieving the highest scores in both BR (50.0%) and OR (40.33%), thereby demonstrating a consistent advantage over competing approaches.
5.4 Ablation Study
@@
Figure 7: Optical Flow Completion. Flows in first row are estimated by RAFT [24]. The completed flows provide motion guidance aligned with the mask surroundings.
Figure 8: Analysis of flow adapter. OR prompts are ”Green leaves of the bush sway in the wind.” and ”A huge rock with flowing water in the background.” respectively.
Effect of flow guidance.
We present the completed flow for the BR and OR tasks, as shown in Fig. 7. The corrupted flow is contextually inpainted with appropriate content that aligns with the surrounding environment while keeping the out-of-mask region intact, which serves as motion guidance for the primary inpainting branch. In Fig. 8, we further demonstrate the effects of flow guidance through two object removal examples. We compare two different variants of our model ($3^{th}$ and $5^{th}$ rows of Tab. 2) to illustrate the impact of flow guidance on inpainting. The inpainted leaves with the flow adapter are more coherent with the surrounding environment (left), especially in the regions at the edges of the object mask. Additionally, the repaired rocks exhibit greater temporal consistency (right) compared to outcomes without the flow adapter which demonstrates flow adapter ensures flow guidance effectively assists the model in generating content consistent with the environment, thereby enhancing temporal coherence and overall quality. Furthermore, as demonstrated in $3^{th}$ and $5^{th}$ rows of Tab. 2, the flow adapter significantly improves FloED’s performance across all metrics, providing substantial benefits to video quality and temporal consistency.
Effect of anchor frame.
As illustrated in Fig. 9, we investigate the impact of the anchor frame. Compared to the variant without the anchor frame strategy, incorporating an anchor frame introduces finer details, thereby enhancing the performance of FloED as an auxiliary strategy. Compared to the flow adapter, its effectiveness is somewhat inferior. As shown in $3^{th}$ and $4^{th}$ rows of Tab. 2, the variant incorporating the flow adapter achieves higher results than the one trained with anchor frame strategy.
Figure 9: Analysis of anchor frame. BR prompt is ”Grilled chicken wings on the barbecue, 8K. ”
@@
Flow Branch L.I. Flow Cache Time SSIM↑
15.55 0.843
$1^{st}$ step 19.79 0.945
$1^{st}$ step $2^{nd}\sim 6^{th}$ 17.72 (↓ 10.5%) 0.944
$1^{st}$ step $2^{nd}\sim 6^{th}$ $6^{nd}\sim 25^{th}$ 16.06 (↓ 18.8%) 0.944
Table 3: Efficiency analysis. L.I. represents the Latent Interpolation. The second row represents no flow-related computation involved. As the closest counterpart, CoCoCo [38] takes 26.25s.
Efficiency analysis.
We conducted an efficiency analysis using a video composed of 16 frames with an A800 GPU. The denoising process is executed for 25 steps with the classifier-free guidance scale set to 15. As presented in Tab. 3, we compare different efficiency strategies against the variant without any flow-related module. As previously mentioned, since our flow branch is independent of time-step during training, during the testing phase, we only need to utilize the flow branch in the first step of the denoising process to complete the damaged optical flow and cache the memory bank. For the remaining denoising steps, we can directly use the completed flow for latent interpolation and the cached K, V for flow guidance. As discussed in Sec. 4.4, we apply latent interpolation exclusively during the initial phase. Fig. 10 illustrates that continuously increasing the acceleration step beyond the early stage of the denoising process results in a sharp decline in performance. By utilizing flow-guided latent interpolation to accelerate the initial denoising and incorporating flow caching in subsequent phases, we can minimize denoising time with only a slight compromise in performance.
Thus, we determined an optimal solution: applying latent interpolation for the initial 5 steps ($2^{nd}\sim 6^{th}$) and flow caching for the remaining steps, resulting in 18.8% speed-up. Compared to the pure variant, which does not incorporate flow completion and flow attention, these efficiency benefits nearly offset the additional computational burden, incurring minimal cost. Notably, FloED outperforms its closest counterpart CoCoCo [38] by 10.19s in total.
@@ Discussion.
Since no trainable modules are involved, latent interpolation can be directly adapted to diffusion-based methods like CoCoCo [38] for acceleration. We acknowledge that obtaining the completed optical flow beforehand may limit its transferability. Additionally, FloED primarily targets BR and OR tasks but is also capable of video editing. Meanwhile, compared to the conventional flow-guided method Propainter [37], FloED demonstrates superior performance in OR during flow completion. These points are further discussed in the appendix.
@@ 6 Conclusion
In this paper, we introduce FloED, an advanced video inpainting framework that efficiently incorporates flow guidance into diffusion models. By employing a dual-branch architecture, FloED first completes corrupted flow through a dedicated flow branch, followed by a multi-scale flow adapter that provides effective motion guidance to the primary inpainting branch. Furthermore, we developed a training-free latent interpolation technique that accelerates the multi-step denoising process. Complemented by a flow attention cache mechanism, FloED minimizes the additional computational burden introduced by incorporating optical flow. Comprehensive experiments demonstrated that FloED outperformed state-of-the-art methods in both background restoration and object removal tasks, showcasing its superior ability to maintain temporal consistency and content coherence in video inpainting.

### v3

@@ 5 Experiments
BR OR
PSNR ↑ SSIM ↑ VFID ↓ ${E_{warp}}$ TC ↑ TA ↑
VC 22.81 0.8614 0.193 3.43 0.987 21.30
CoCoCo 23.08 0.8694 0.165 3.73 0.991 21.97
DiffuEraser 24.23 0.8583 0.218 2.98 0.984 19.19
Ours 29.17 0.9441 0.118 2.83 0.994 22.49
Table 1: Quantitative comparisons. Metrics including PSNR, SSIM, VFID, Optical flow warping error(${E_{warp}}$), Temporal consistency (TC), and Text alignment (TA).
@@
Dataset and Benchmark.
We utilize the Open-Sora-Plan dataset [13], splitting videos at scene cuts to obtain 421,396 high-quality video clips paired with captions. We further developed an evaluation benchmark comprising 100 previously unseen videos sourced from Pexels and Pixabay platforms, with 50 designated for object removal (OR) and 50 for background restoration (BR). For BR task, we use synthetic random masks that focus on the background. For the OR task, object masks are obtained by applying Segment-Anything [11] (SAM) to each frame. Each video is manually selected to ensure a diverse range of motion amplitudes and camera movement speeds, while guaranteeing 4K resolution and 100 frames in total. For the caption, we generate the corresponding video prompts using VideoGPT [16], while manually revising appropriate background prompts for OR.
Training and Inference Details.
We employed a two stage training strategy with a resolution of 512 and 16-frame video sequences. And we generated mask sequences with random directions, and shapes to simulate BR and OR tasks. The first stage is trained on 8 NVIDIA A800 GPUs for 5 epochs with a batch size of 8. And second stage is trained on 8 NVIDIA A800 GPUs for 30 epochs with a batch size of 128, which is achieved through gradient accumulation the $\lambda$ is set to 0.1 during the second stage. During inference, we use DDIM [9] and empirically define the speed-up step $S$ as 5 (25 steps in total).
@@ 5.2 Comparisons
We present comprehensive comparisons with open-sourced text-guided diffusion-based approaches, including VideoComposer [24], CoCoCo [36] and DiffuEraser [14].
@@
As demonstrated in Fig. 3, VideoComposer, CoCoCo and DiffuEraser exhibit persistent limitations especially in OR tasks, frequently generating visual artifacts and content hallucinations that disrupt semantic consistency with scene context. In contrast, FloED inpaints the mask region with compatible contents and demonstrates precise text-conditioned generation capabilities, achieving superior temporal consistency and overall coherence in both BR and OR tasks.
Figure 4: Optical flow related ablation studies. (F) ablation study demonstrates FloED conducts flow warping at noise $\epsilon$ instead of $\mathbf{z}_{0}$.
@@
Figure 5: We conduct reliable User Study with randomized order to assess inpainting outcomes of 4 methods.
Quantitative Comparisons.
We take metric evaluation and user studies to demonstrate quantitative comparisons.
(1) Metric evaluation. For BR tasks, we employ the PSNR [26], VFID [23], and SSIM [25] to quantify basic quality. Additionally, we assess temporal consistency using flow warping error [5, 12] in conjunction with Temporal Consistency (TC) [33]. TC is measured by the cosine similarity between consecutive frames in the CLIP-Image [8] feature space. For OM tasks, since ground truth data is unavailable for evaluating the aforementioned metrics, we utilize Text Alignment (TA) [8] as an evaluation metric, which also leverages the CLIP score. For consistency, all metric evaluations were conducted at a resolution of 512×512. As shown in Tab. 1, FloED outperforms other methods in all the metrics, which demonstrates state-of-the-art performance.
(2) User study. Since CLIP scores do not always align with human perception [33], we conducted a comprehensive user study in which 15 annotators evaluated the inpainting results across both BR and OR tasks (100 videos), assessing temporal consistency, text alignment, and context compatibility to select the best one among 4 methods.As illustrated in Fig. 5, our model was highly favored, achieving the highest scores in both BR (62.27%) and OR (56.40%).
5.3 Ablation Studies
Flow-related Ablation Studies.
We conduct flow-related ablation studies to validate the motion guidance, with experimental results shown in Fig. 4.
(1) Flow completion. Using an object removal scenario as a case study, completed optical flow results demonstrate that the corrupted flow undergoes context-aware inpainting with spatially and temporally coherent content which maintains alignment with the surrounding environment (comparing B with C). And the reconstructed results further validate the completion capability of our time-agnostic flow branch.
(2) Flow adapter. These reconstructed flows provide crucial motion guidance to the primary inpainting branch, effectively enhancing context compatibility and achieving higher video coherence through multi-scale flow adapters. Comparing the outcomes of D with E, black artifacts and incompatible lighting condition can be largely reduced. The experimental findings demonstrate that the multi-scale flow adapters’ injection of motion guidance significantly improves environmental consistency in generated content, which in turn enhances temporal coherence and overall quality. Meanwhile, we also conduct the quantitative architecture ablation study in Tab. 2. Compared with anchor frame strategy, the multi-scale flow adapters demonstrate superior efficacy in enhancing framework’s performance, confirming their critical role in FloED. More ablation study results are provided in the Appendix.
(3) Flow warping operation. Instructively, our ablation study F reveals that applying flow warping to the intermediate noise estimate $\epsilon$ during latent interpolation leads to cumulative error propagation with blurriness results compared to baseline D. To circumvent this, we reformulate the warping operation in the clean latent space $\mathbf{z}_{0}$(as per Equation 4).
@@ 5.3 Ablation Studies
Flow Branch Latent Interpolation Flow Cache Average time per frame(s)
0.1287
$1^{st}$ step 0.1491
$1^{st}$ step $2^{nd}\sim 6^{th}$ 0.1342 (↓ 9.9%)
$1^{st}$ step $2^{nd}\sim 6^{th}$ $6^{nd}\sim 25^{th}$ 0.1291 (↓ 13.4%)
Table 3: Efficiency ablation study (432 $\times$ 240). The second row represents variant with no flow-related module involved (blue).
Efficiency.
In this section, we conduct the efficiency experiments with single H800 NVIDIA GPU under FP16 setting. And FloED’s denoising process is executed for 25 steps with the classifier-free guidance scale (CFG) $>$ 1.
(1) Latent interpolation steps. As discussed in Sec. 4.4, we apply latent interpolation exclusively during the initial phase. Thus, we conduct a speeding steps study in Fig. 6. As depicted in Fig. 6, continuously increasing the acceleration step results in a sharp decline in performance when the speeding steps $S$ beyond the early stage of the denoising process. Our experimental findings demonstrate that utilizing flow-guided latent interpolation for the initial 5 steps, we can minimize denoising time with only a slight compromise in performance.
(2) Efficiency ablation study. As presented in Tab. 3, we compare different efficiency strategies against the variant without any flow-related module. As previously mentioned, since our flow branch is independent of time-step during training, during the testing phase, we only need to utilize the flow branch in the first step of the denoising process to complete the damaged optical flow and cache the memory bank. For the remaining denoising steps, we can directly use the completed flow for latent interpolation and the cached K, V for flow guidance.
Thus, we determined an optimal solution: applying latent interpolation for the initial 5 steps ($2^{nd}\sim 6^{th}$) and leveraging flow caching for the remaining steps for complement, resulting in 13.4% speed-up in the resolution of 432$\times$240. Compared to the pure variant, which does not incorporate flow completion and flow attention, these efficiency benefits nearly offset the additional computational burden, incurring minimal cost.
(3) Efficiency comparisons. As shown the Tab. 4, under the same denoising steps, FloED outperform other diffusion-based counterparts, such as CoCoCo and DiffuEraser, across all resolutions, demonstrating its state-of-the-art efficiency.
Average time per frame (s) 512$\times$512 432$\times$256 256$\times$256
Ours 0.3412 0.1291 0.0756
CoCoCo 0.4033 0.2084 0.1735
DiffuEraser 0.4176 0.5751 1.0336
Table 4: Efficiency comparisons (CFG $>$ 1, 25 steps, FP16).
@@ Discussion.
This paper focuses on text-guided video inpainting, primarily comparing with diffusion-based solutions. Our FloED also demonstrates superior performance over conventional flow-guided methods like ProPainter [34], with detailed comparisons provided in the Appendix. Furthermore, FloED’s latent interpolation can be directly extended to other diffusion-based approaches (e.g., CoCoCo [36]) for accelerated processing. However, we note that the strategy of pre-completing corrupted optical flow might restrict its transferability across different application scenarios.
@@ 6 Conclusion
In this paper, we introduced FloED, a coherent video inpainting framework that effectively integrates optical flow guidance into diffusion models to enhance temporal consistency and computational efficiency. By employing a dual-branch architecture, FloED first reconstructs corrupted flow, which then guides the inpainting process through multi-scale flow adapters. Additionally, our training-free latent interpolation technique and flow attention cache significantly reduce the computational overhead typically associated with optical flow integration. Experimental results demonstrate that FloED achieves state-of-the-art performance in both background restoration and object removal, showcasing its superior ability to maintain temporal consistency and content coherence in video inpainting.